OpenStudy (anonymous):

solution again! 2x-y=-7 4x-y=-4

hartnn (hartnn):

Subtract those 2 equations, what you get ?

OpenStudy (shaik0124):

multiply first equation by 2 and subtract both equations

OpenStudy (anonymous):

y=18?

OpenStudy (anonymous):

y=10 and x=3/2

OpenStudy (anonymous):

use substitution method

OpenStudy (anonymous):

can you show me how you got that? this stuff confuses me. seem to always to it wrong

OpenStudy (anonymous):

take first equation 2x-y=-7

OpenStudy (anonymous):

k

OpenStudy (anonymous):

2x=-7+y y=2x+7.......(1) put value of y in second equation

OpenStudy (anonymous):

ll be 4x-(7+2x)=-4

OpenStudy (anonymous):

4x-7-2x=-4

OpenStudy (anonymous):

2x=7-4

OpenStudy (anonymous):

x=3/2